OpenStudy (anonymous):

Simplify the expression. [3x^4y/xy^-2]^-2

OpenStudy (anonymous):

\[x^a * x^b = x^{a+b}\]\[\frac{ x^a }{ x^b } = x^{a - b}\]\[(x^a)^b = x^{ab}\]

OpenStudy (anonymous):

we have:\[\left( \frac{ 3x^4y }{ xy^{-2} } \right)^{-2}\] inside the bracket first:\[\left( 3x^{4-1}y^{1- (-2)} \right)^{-2} = \left( 3x^3y^3 \right)^{-2} = 3^{-2}x^{3(-2)}y^{3(-2)} = \frac{ 1 }{ 9 }x^{-6}y^{-6} = \frac{ 1 }{ 9x^6y^6 }\]

OpenStudy (anonymous):

thank you Euler271!

OpenStudy (anonymous):

glad i could help :)